OpenStudy (anonymous):

A lion accelerates from rest at a rate of 4.2 m/s2. How far has the lion gone after 4 seconds?

OpenStudy (anonymous):

\[x _{f}=x _{i}+v _{i}\Delta t + 1/2a(\Delta t)^{2}\] \[x _{f}= 0 + (0)(4) + 1/2(4.2)(4)^{2}\] \[x _{f}=33.6\]

OpenStudy (anonymous):

thanks

mathslover (mathslover):

@larquitta1995 - Welcome to OpenStudy first of all. \(\Huge{ \boxed{\boxed{\LARGE{\stackrel{\color{red}{\ast\qquad\star\qquad\ast\qquad\star\qquad\ast\qquad\star\qquad\ast\qquad\star}}{\mathscr {\color{orange}{Welcome } \quad \color{green}{To}\quad \color{purple}{OpenStudy }!!}}}\\{\color{pink}{\ast\qquad\star\qquad\ast\qquad\star\qquad\ast\qquad\star\qquad\ast\qquad\star}}}}}\) If you think @RoboCozz helped you in your doubt then do spare some time to give him a medal (best response) which shows that he had helped you! Thanks!